OpenStudy (anonymous):

22/7 + 11/14 = whats the quotient?

OpenStudy (anonymous):

55/14

ganeshie8 (ganeshie8):

\(\huge \frac{22}{7} + \frac{11}{14}\) \(\huge \frac{3*7+1}{7} + \frac{11}{14}\) \(\huge 3 \frac{1}{7} + \frac{11}{14}\) \(\huge 3 + \frac{1}{7} + \frac{11}{14}\) \(\huge 3 + \frac{1*2}{7*2} + \frac{11}{14}\) \(\huge 3 + \frac{2}{14} + \frac{11}{14}\) \(\huge 3 + \frac{2+11}{14} \) \(\huge 3 + \frac{13}{14} \) \(\huge 3 \frac{13}{14} \) \(\huge \color{purple}{ => quotient = 3}\)